OpenStudy (anonymous):

12x=9x+7 Explain how to do to this step by step /).(\ I need a refresher.. I'm new to this app soo I guess people are giving metals to people who help them.. So I'll do that too ☺️💕

pooja195 (pooja195):

Hi Welcome to openstudy :-) How would you approach this problem? Where would you start?

OpenStudy (anonymous):

No I don't think so.. It just says solve in the math book XD

OpenStudy (anonymous):

I think you would start by adding both X's together?

pooja195 (pooja195):

Correct but let's subtract instead :) 12-9=?

OpenStudy (anonymous):

Or wait.. Subtract the 9x from 12x and the rest is your answer?

pooja195 (pooja195):

Nope there's more what is 12-9?

OpenStudy (anonymous):

3

pooja195 (pooja195):

Good now we have 3x=7 any ideas on what should be done next?

OpenStudy (anonymous):

Divide the 3x by 7

pooja195 (pooja195):

Yes :) 3x/3=7/3 solve for x

pooja195 (pooja195):

Don't over think it

OpenStudy (anonymous):

Thank you ☺️💕🏅🏅🏅🏅🏅

pooja195 (pooja195):

^_^ yw ^_^ <3